Consider a sample with data values of 19, 20, 5 ,8 , and 18 .

Compute the variance.

(to 1 decimal)

Compute the standard deviation.

(to 2 decimals)
Answer #1

Solution :

n = 5

\bar x = \sum x / n = (19 + 20 + 5 + 8+18) / 5 = 70 / 5 = 14

s2 =  \sum(x - \bar x)2 / n - 1  

= (19 - 14)2 + (20 - 14)2 + (5 - 14)2 + (8 - 14)2 + (18 - 14)2 / 5 - 1

= 194 / 4

= 48.5

variance = 48.5

standard deviation = \sqrtvariance = \sqrt48.5 = 6.96